OpenStudy (anonymous):

2x – 3y = –10, when x = 4, What is the value of y

jimthompson5910 (jim_thompson5910):

replace x with 4 to get 2x – 3y = –10 2*4 – 3y = –10 8 – 3y = –10 now solve for y

OpenStudy (anonymous):

For 2x – 3y = –10, when x = 4, What is the value of y? negative two-thirds two-thirds 6 –6

OpenStudy (anonymous):

wich one is it